# Quantum effects and temperature dependence of the free energy barrier and shear frequency in bilayer graphene

### v2

@@ Abstract
Little is known about the quantum effects of friction, and first-principles studies are virtually non-existent. In multilayer graphene, the order of the stacking of the layers plays a crucial role in the electronic properties and the manifestation of superconductivity, and friction affects how layers slide to change the stacking. Here, we study the effects of friction in bilayer graphene by focusing on quantum fluctuations in the free energy barrier that layers have to overcome to slide with respect to each other. We evaluate the free energy at the barrier in bilayer graphene through the self-consistent harmonic approximation, accounting for quantum and thermal ionic fluctuations, and obtain a very large reduction of the barrier of more than 30% at low temperatures. There is little temperature dependence up to about 500 K, but it increases at higher temperatures. Our approach is general and paves the way for systematically accounting for nuclear quantum and thermal effects in free energy barriers of other macroscopic systems.
@@ I Introduction
There has been intense research on graphene since its discovery in 2004, and also on multilayer graphene systems. Among many other interesting properties, flat bands are present in multilayer rhombohedral (ABC-stacked) graphene (RG)Koshino (2010), and in fact, superconductivity has been observed in RG in 2021Zhou et al. (2021). Several works recently studied, both experimentally and theoretically, under which conditions RG can form Yankowitz et al. (2014); Geisenhof et al. (2019); Bouhafs et al. (2020); Gao et al. (2020); Li et al. (2020); Kerelsky et al. (2021). In particular, some of us unveiled how the stacking of graphene layers changes under shear stress Nery et al. (2020), and obtained that it produces many consecutive layers of RG. In fact, all samples with long-range RG that have been observed so far have been subject to some form of shear stress Lin et al. (2012); Henni et al. (2016); Henck et al. (2018); Yang et al. (2019); Shi et al. (2020); Bouhafs et al. (2020), and thus of friction.
@@
To move graphene layers with respect to each other, they have to overcome a free energy barrier that separates different minima. When doing so, they experiment “stick-slip” motion: they move gradually with increasing shear stress, followed by sudden rearrangements and decrease of the stressLiu et al. (2012). The maximum shear-stress to “unlock” layers is correlated with the height of the barrier. Such barrier corresponds to the activation energy of a metastable state, where a graphene flake (from nanometers up to possibly micrometer scales) sitting on top of a large graphene layer is displaced relative to the equilibrium position. In clean samples with micrometer sized flakes however, different regions can have different stackings (AB and AC)Henck et al. (2018); Jiang et al. (2018), and a domain wall in the metastable state separates both regions. The displacement of layers in this case is more complex and is not considered in our work.
@@
Superlubricity is another phenomenon present in systems involving graphene layers or graphite that has attracted a lot of attention(Liu et al., 2014; Sinclair et al., 2018), in which the energy landscape corresponding to the relative displacement of graphene layers is also central. Yet another phenomena involving sliding surfaces is thermolubricity, which refers to the assistance from thermal excitations to overcome barriers. This term has been coined by some works that have studied the transition from stick-slip motion into stochastic fluctuations in friction at sufficiently high temperaturesBaykara et al. (2018). A few works have studied how quantum effects, including zero-point renormalization (ZPR) and tunneling, affect the free-energy barrier of dissociative adsorptionMills and Jónsson (1994) and diffusionPoma et al. (2012); McIntosh et al. (2013); Cendagorta et al. (2016) of hydrogen. However, there are no studies of the quantum effects and temperature dependence of free-energy barriers in macroscopic systems.
@@
Another aspect of multilayer graphene and graphite that still needs further investigation and that serves as a consistency check on the adequacy of our approach is thermal expansion (TE). Ref. McQuade et al., 2021 measured it in bilayer and trilayer graphene for the first time. Ref. Mittal et al., 2021 uses the quasi-harmonic approxmimation (QHA), which underestimates the out-of-plane TE at low temperaturesBailey and Yates (1970) and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperaturesBailey and Yates (1970) and overestimates it at high temperaturesMorgan (1972). Much better out-of-plane results were obtained by using path integral molecular dynamics (PIMD)Herrero and Ramírez (2020), but while the TE flattens at 500 K in graphite, experimental values continue increasingMorgan (1972); Marsden et al. (2018). However, in-plane PIMD results of Ref. Herrero and Ramírez, 2020 vastly overestimate the TE.
Motivated by the aforementioned phenomena, especially by the lack of understanding of quantum effects in friction, we focus on studying the ZPR and temperature effects of the free-energy barrier. This is a very challenging problem due to the difficulty of determining the free energy in a non-equilibrium position. Here we tackle it using the state-of-the-art stochastic self-consistent harmonic approximation (SSCHA)Monacelli et al. (2021). First, in Sec.II, we briefly introduce SSCHA and the QHA, and explain how is the barrier determined. Then, in Sec. III, we present our main results and compare to the height of the barrier with fixed atoms. Subsequently, using also the QHA, we calculate the temperature dependence of the shear frequency and compare to the limited experimental data available. Finally, we determine the in-plane and out-of-plane TE coefficient in bilayer graphene and graphite, and compare our results to previous theoretical and experimental work. In Sec. IV, we summarize the conclusions of our work.
@@ II Theoretical and Computational Framework
Here, we briefly describe the QHA, SSCHA, and a new SSCHA interpolation method used to determine the barrier. We also introduce the equation that is used to fit SSCHA thermal expansion results. A more detailed description of SSCHA can be found in Refs. Errea et al., 2014 or Monacelli et al., 2021.
@@ II.1 Quasi-harmonic approximation (QHA)
In the QHA, at each value of the lattice parameters $\{a_{i}\}$, the free energy $F(T,\{a_{i}\})=U(\{a_{i}\})-TS(T,\{a_{i}\})$ is given by the standard harmonic expression. That is,
$\begin{split}F(T,\{a_{i}\})=&U_{\mathrm{lat}}(\{a_{i}\})+\frac{1}{N}\sum_{\mathbf{q}\nu}\frac{1}{2}\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})+\\ &+\frac{1}{N}\sum_{\mathbf{q}\nu}kT\mathrm{ln}\left(1-e^{-\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})/kT}\right).\\ \end{split}$ (1)
The value of the lattice parameters at a given temperature, $a_{i}(T)$, is determined by minimizing the free energy. Using an interatomic potential, the phonon frequencies at each value of the lattice parameters can be determined with little computational effort, and thus the QHA offers a quick method to determine the lattice expansion. Using first-principles calculations, phonons can be determined using density functional perturbation theory (DFPT), which uses a primitive cell and is much faster than frozen phonon (finite differences) calculations in supercells. However, it does not account for the “true anharmonicity”Allen (2019), present at fixed values of the parameters when varying temperature. For example, it gives an increasing as opposed to decreasing temperature dependence of the G mode in graphene, due to the negative TE of the lattice parameter. Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see later Fig. A10), then an acoustic mode becomes unstable and the QHA is ill-defined. These issues are not present in SSCHA.
@@ II.2 Stochastic self-consistent harmonic approximation (SSCHA)
where $\mathbf{R}$ indicates a general ionic configuration and $\rho_{\mathcal{H}}(\mathbf{R})=\langle\mathbf{R}|e^{-\beta\mathcal{H}}|\mathbf{R}\rangle/Z_{\mathcal{H}}$ is the probability density of finding the crystal in a generic configuration $\mathbf{R}$ ($Z_{\mathcal{H}}=\mathrm{tr}[e^{-\beta\mathcal{H}}]$ is the partition function). By stochastically sampling configurations $\mathbf{R}$ for a given density matrix $\rho_{\mathcal{H}}$, and determining the gradient of the free energy with respect the harmonic parameters (centroids and force constant matrix elements), the free energy can be minimized self-consistently. Thanks to the variational ansatz, the SSCHA quantum free energy is rigorously defined, even in out-of-equilibrium and unstable configurations like SP. Notably, on the other hand, the QHA cannot determine the barrier, since the harmonic frequencies are not well defined at SP.
The logarithm term in the harmonic term Eq. (1) can be hard to converge for low frequencies. Separating the total free energy as in Eq. (4) allows to interpolate the auxiliary harmonic term up to a very fine $q$-mesh, which is fundamental to avoid oversampling the energy of the long-ranged shear modes, while the anharmonic contribution can be properly converged already with small simulation cells. So rather than calculating $F_{\mathcal{H}}$ directly with the grid corresponding to the SC used in the minimization procedure, the force constant matrix can be Fourier interpolated in the usual way, and $F_{\mathcal{H}}$ can be calculated using a dense mesh.
This procedure is followed both at SP and the equilibrium position, and the barrier is obtained by taking the difference of the corresponding free energies. More technical details are provided in the Appendix, Sec. A1.
The SSCHA frequencies are in principle auxiliary frequencies that minimize the free energy. In a purely harmonic system, such frequencies correspond to the physical frequencies. But in general, they have to be corrected with the so-called “bubble correction” (and also an additional higher order term in very anharmonic cases)Monacelli et al. (2021), which we do in Sec. III.2 to obtain the temperature dependence of the shear frequency. Convergence with the supercell size of both the physical and auxiliary frequencies can be seen in the Appendix, Fig. A5.
II.3 Thermal expansion (TE)
As opposed to the barrier, the thermal expansion (TE) can also be calculated using the QHA, which although is much faster, is also in principle less accurate. With both QHA and SSCHA we determine the in-plane and out-of-plane lattice parameters as a function of temperature, $a(T)$ and $c(T)$, to then determine the in-plane and out-of-plane coefficients of TE, $\alpha_{\mathrm{a}}(T)=\tfrac{1}{a}\tfrac{da}{dT}$ and $\alpha_{\mathrm{c}}(T)=\tfrac{1}{c}\tfrac{dc}{dT}$. In the QHA case, the TE can calculated quickly and accurately for many temperatures, so we just use finite differences. For SSCHA, to reduce numerical error, we use Eq. (6) of Ref. Zorzi and Perottoni, 2021,
$b(T)=b_{0}\mathrm{exp}\left(\frac{X\theta}{e^{X/T}-1}\right),$ (5)
which depends on parameters $b_{0}$, $X$ and $\theta$, to fit $a(T)$ and $c_{d}(T)$ ($b$ can be $a$ or $c_{d}$). Then we determine the coefficients of TE from the derivative of the fit.
@@ III.1 Barrier
We first focus on studying bilayer graphene, where the in-plane and (most importantly) out-of-plane interactions are given by the interatomic potential of Ref. Libbi et al., 2020 and Ref. Wen et al., 2018, respectively. The out-of-plane interactions determine the barrier. The energy profile of a layer moving with respect to the other one along the bond direction is given by Fig. 1, which gives good reults when comparing to theory and experiment (see Table 1).
@@
The lower layer is fixed in position A, and there are two minima of the same energy, AB and AC, at a distance of a bond length $d=1.42$. For layers to slide with respect to each other, they have to overcome a barrier $V_{\mathrm{SP}}$. SP is a saddle point in the full energy landscape, so the shear mode in the perpendicular direction (along the direction of two nearby AA maxima) is stable. In order to determine the ZPR and temperature dependence of the barrier, we use SSCHA to determine the free energy at the equilibrium position and SP, using the interpolation method described in Sec. II.2.
@@
Figure 2: Barrier as a function of temperature. It is reduced from 1.24 meV/atom, the value at fixed nuclei (VSP in Fig. 1), to about 0.8 meV/atom, corresponding to a large reduction of about 35% due to anharmonicities. The barrier does not change much up to about 500 K, but the value at 1000 K is about 20% smaller relative to the value at 100 K. The lines are a guide to the eye, as in the other figures of this work.
@@
To confirm that quantum fluctuations are very minor down to 100 K, we performed SSCHA calculations in the classical limit $\hbar\omega/kT\rightarrow 0$ (which recovers the classical Boltzmann distribution for the distorted configurations) at 100 K and 300 K. We obtained comparable results, except for a small difference of less than 0.05 meV/atom at 100 K with respect to the quantum values. Much lower temperatures would have to be considered to recover the classical value at fixed ions. Thus, it is the anharmonic vibrations that play a decisive role in the ability of layers to slide with respect to each other.
In a previous workNery et al. (2020), using density functional theory with an LDA functional (see Table 1), some of us obtained that the maximum shear stress in simulated stick-slip motion is 0.2 GPa, and we mentioned a good agreement with experimental values 0.14 GPa (Ref. Liu et al., 2012) and 0.1 GPa (Ref. Dienwiebel et al., 2004). Assuming the maximum shear stress decreases by 35% just like the barrier, we obtain 0.13 GPa, in excellent agreement with experiments.
Other systems involving hydrogen diffusion across a free energy barrier do present significant differences between classical and quantum free energy barriers due to the large quantum fluctuations of hydrogen. Such differences have been observed at 100 K (Ref. Mills and Jónsson, 1994), 250 K (Ref. McIntosh et al., 2013), and even higher temperaturesPoma et al. (2012), although other worksCendagorta et al. (2016) obtain very similar barriers at already 200 K.
Regarding the evolution of the barrier as the temperature increases, there is a weak dependence up to about 500 K, since the barrier remains close to around 0.8 meV/atom. However, when temperature increases up to 1000 K, the barrier is further reduced by 20% relative to the value at 100 K. Thus, high temperatures should help to reduce (static) friction between layers. In particular, it could aid the production of ABC-stacked multilayer graphene from its more common AB-stacked version through shear stress, as mentioned in the Introduction. More in general, high temperatures could also help engineer specific stacking arrangements in other layered materials.

### v3

# Free energy barrier and thermal-quantum behavior of sliding bilayer graphene

@@ Abstract
In multilayer graphene, the stacking order of the layers plays a crucial role in the electronic properties and the manifestation of superconductivity. By applying shear stress, it is possible to induce sliding between different layers, altering the stacking order. Here, focusing on bilayer graphene, we analyze how ionic fluctuations alter the free energy barrier between different stacking equilibria. We calculate the free energy barrier through the state-of-the-art self-consistent harmonic approximation, which can be evaluated at unstable configurations. We find that above 100 K there is a large reduction of the barrier of more than 30% due to thermal vibrations, which significantly improves the agreement between previous first-principles theoretical work and experiments in a single graphite crystal. As the temperature increases, the barrier remains nearly constant up to around 500 K, with a more pronounced decrease only at higher temperatures. Our approach is general and paves the way for systematically accounting for thermal effects in free energy barriers of other macroscopic systems.
@@ I Introduction
There has been intense research on graphene since its discovery in 2004, and also on multilayer graphene systems. Among many other interesting properties, flat bands are present in multilayer rhombohedral (ABC-stacked) graphene (RG)Koshino (2010), and superconductivity has been observed in RG in 2021Zhou et al. (2021). Several works recently studied, both experimentally and theoretically, under which conditions RG can form Yankowitz et al. (2014); Geisenhof et al. (2019); Bouhafs et al. (2021); Gao et al. (2020); Li et al. (2020); Kerelsky et al. (2021). In particular, some of us unveiled how the stacking of graphene layers changes under shear stress Nery et al. (2020), and obtained that it produces many consecutive layers of RG. In fact, all samples with long-range RG that have been observed so far have been subject to some form of shear stress Lin et al. (2012); Henni et al. (2016); Henck et al. (2018); Yang et al. (2019); Shi et al. (2020); Bouhafs et al. (2021).
@@
To move graphene layers with respect to each other, they have to overcome a free energy barrier that separates different minima. When doing so, they experiment “stick-slip” motion: they move gradually with increasing shear stress, followed by sudden rearrangements and decrease of the stressLiu et al. (2012). The maximum shear-stress to “unlock” layers is correlated with the height of the barrier. Such barrier corresponds to the activation energy of a metastable state, where a graphene flake sitting on top of a large graphene layer is displaced relative to the equilibrium position.
@@
Superlubricity is a another phenomenon where the energy landscape corresponding to the relative displacement of layers is also central. In this mechanism, which has attracted a lot attention(Liu et al., 2014; Sinclair et al., 2018), particularly in carbon-based systems like multilayer graphene and graphite, incommensurate surfaces (i.e. atomic lattices without matching periodicity) lead to extremely low friction. Yet another phenomena involving sliding surfaces is thermolubricity, which refers to the assistance from thermal excitations to overcome barriers. This term has been coined by some works that have studied the transition from stick-slip motion into stochastic fluctuations in friction at sufficiently high temperaturesBaykara et al. (2018). Despite the interest in sliding surfaces, there are no studies that account for both temperature and quantum effects of free-energy barriers in macroscopic systems. A few works have studied quantum effects in free-energy barriers, including zero-point renormalization and tunneling, but of systems involving hydrogen molecules, including dissociative adsorptionMills and Jónsson (1994) and diffusionPoma et al. (2012); McIntosh et al. (2013); Cendagorta et al. (2016).
The variation of the potential along a periodic crystal surface, combined with the attraction between layers, prevents free sliding between them. This phenomenon has been extensively studied in various systems, like graphiteDienwiebel et al. (2004), MoS2Li et al. (2016), silicon on goldLiu et al. (2015), among others. When using a tip with the goal of sliding layers, normal forces need not be large. For pressures much lower than 1GPa, the alteration of the free-energy landscape is minimal (see SI of Ref. Nery et al., 2020), so the relaxed barrier remains the dominant contribution that impedes sliding. In addition, some works on superlubricity have observed that friction has an extremely weak dependence with the load, attributed for exampleDienwiebel et al. (2004) to a contact area that remains nearly constant in the force range applied. Here, we focus on calculating the height of the barrier, which should be approximately proportional to the force –given by the maximum slope of the free energy profile– that is needed for a layer to move from one minima to another in commensurate systems.
@@
Another aspect of multilayer graphene and graphite that still needs further investigation and that serves as a consistency check on the adequacy of our approach is thermal expansion (TE). Ref. McQuade et al., 2021 measured it in bilayer and trilayer graphene for the first time. Ref. Mittal et al., 2021 uses the quasi-harmonic approximation (QHA), which underestimates the out-of-plane TE at low temperaturesBailey and Yates (1970) and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperaturesBailey and Yates (1970) and overestimates it at high temperaturesMorgan (1972). Much better out-of-plane results were obtained by using path integral molecular dynamics (PIMD)Herrero and Ramírez (2020), but while the TE flattens at 500 K in graphite, experimental values continue increasingMorgan (1972); Marsden et al. (2018). However, in-plane PIMD results of Ref. Herrero and Ramírez, 2020 vastly overestimate the TE.
Motivated by the aforementioned phenomena, especially by the possibility of facilitating sliding through temperature, we focus on thermal effects on the free-energy barrier. This is a very challenging problem due to the difficulty of determining the free energy in a non-equilibrium position. Here we tackle it using the state-of-the-art stochastic self-consistent harmonic approximation (SSCHA)Errea et al. (2014); Monacelli et al. (2021). First, in Sec. II, we briefly introduce SSCHA and the QHA, and explain how the barrier is determined. Then, in Sec. III, we present our main results and compare to the height of the barrier with fixed atoms. Subsequently, using also the QHA, we calculate the temperature dependence of the shear frequency and compare to the limited experimental data available. Finally, we determine the in-plane and out-of-plane TE coefficient in bilayer graphene and graphite, and compare our results to previous theoretical and experimental work. In Sec. IV, we summarize the conclusions of our work.
@@ II Theoretical and Computational Framework
Here, we briefly describe the QHA, SSCHA, and a new SSCHA interpolation method used to determine the barrier. A more detailed description of SSCHA can be found in Refs. Errea et al., 2014 or Monacelli et al., 2021.
@@ II.1 Quasi-harmonic approximation (QHA)
In the QHA, at each value of the lattice parameters $\{a_{i}\}$, the free energy $F(T,\{a_{i}\})=U(\{a_{i}\})-TS(T,\{a_{i}\})$, where $U$ is the internal energy, $S$ the entropy and $T$ the temperature, is given by the standard harmonic expression. That is,
$\begin{split}F(T,\{a_{i}\})=&U_{\mathrm{lat}}(\{a_{i}\})+\frac{1}{N_{q}}\sum_{\mathbf{q}\nu}\frac{1}{2}\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})+\\ &+\frac{1}{N_{q}}\sum_{\mathbf{q}\nu}kT\mathrm{ln}\left(1-e^{-\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})/kT}\right).\\ \end{split}$ (1)
where $U_{\mathrm{lat}}\{a_{i}\}$ is the internal energy of the lattice for a given value of the lattice parameters, $N_{q}$ the number of phonon wavevectors $\mathbf{q}$, $\nu$ the mode, and $\omega_{\mathbf{q}\nu}$ the corresponding phonon frequency.
The value of the lattice parameters at a given temperature, $a_{i}(T)$, is determined by minimizing the free energy. Using an interatomic potential, the phonon frequencies at each value of the lattice parameters can be determined with little computational effort, and thus the QHA offers a quick method to determine the lattice expansion. Using first-principles calculations, phonons can be determined using density functional perturbation theory, which uses a primitive cell and is much faster than frozen phonon (finite differences) calculations in supercells. However, it does not account for the “true anharmonicity”Allen (2019), present at fixed values of the parameters when varying temperature. For example, it gives an increasing as opposed to decreasing temperature dependence of the G mode in graphene, due to the negative TE of the lattice parameter. Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see the Supplementary Information (SI), Fig. S11), then an acoustic mode becomes unstable and the QHA is ill-defined. These issues are not present in SSCHA.
@@ II.2 Stochastic self-consistent harmonic approximation (SSCHA)
where $\mathbf{R}$ indicates a general ionic configuration and $\rho_{\mathcal{H}}(\mathbf{R})=\langle\mathbf{R}|e^{-\beta\mathcal{H}}|\mathbf{R}\rangle/Z_{\mathcal{H}}$ is the probability density of finding the crystal in a generic configuration $\mathbf{R}$ ($Z_{\mathcal{H}}=\mathrm{tr}[e^{-\beta\mathcal{H}}]$ is the partition function). By stochastically sampling configurations $\mathbf{R}$ for a given density matrix $\rho_{\mathcal{H}}$, and determining the gradient of the free energy with respect the harmonic parameters (centroids and force constant matrix elements), the free energy can be minimized self-consistently. SSCHA is thus a nonperturbative method, as it does not rely on a perturbative expansion: it explores configurations with amplitudes that depend on temperature, without being limited to small displacements.
Thanks to the variational ansatz, the SSCHA quantum free energy is rigorously defined, even in out-of-equilibrium and unstable configurations like SP. In contrast, the QHA cannot determine the barrier, since the harmonic frequencies are not well defined at SP.
The logarithm term in the harmonic term Eq. (1) can be hard to converge for low frequencies. Separating the total free energy as in Eq. (4) allows to interpolate the auxiliary harmonic term up to a very fine $q$-mesh, which is fundamental to avoid oversampling the energy of the long-ranged shear modes, while the anharmonic contribution can be properly converged already with small simulation cells. So rather than calculating $F_{\mathcal{H}}$ directly with the grid corresponding to the SC used in the minimization procedure, the force constant matrix can be Fourier interpolated in the usual way (building the force constant in real space from the existing $q$-grid, and then Fourier transforming back to any desired value of $\mathbf{q}$), and $F_{\mathcal{H}}$ can be calculated using a dense mesh.
This procedure is followed both at SP and the equilibrium position, and the barrier is obtained by taking the difference of the corresponding free energies. More technical details are provided in the SI, Sec. S1.
The SSCHA frequencies are in principle auxiliary frequencies that minimize the free energy. In a purely harmonic system, such frequencies correspond to the physical frequencies. But in general, they have to be corrected with the so-called “bubble correction” (and also an additional higher order term in very anharmonic cases)Monacelli et al. (2021), which we do in Sec. III.2 to obtain the temperature dependence of the shear frequency. Convergence with the supercell size of both the physical and auxiliary frequencies can be seen in the SI, Fig. S6.
@@ III.1 Barrier
We first focus on studying bilayer graphene, where the in-plane and (most importantly) out-of-plane interactions are given by the interatomic potential of Ref. Libbi et al., 2020 and Ref. Wen et al., 2018, respectively. The out-of-plane interactions determine the barrier. The energy profile of a layer moving with respect to the other one along the bond direction is given by Fig. 1, which gives good results when comparing to theory and experiment (see Table 1).
@@
The lower layer is fixed in position A, and there are two minima of the same energy, AB and AC, at a distance of a bond length $d=1.42$. For layers to slide with respect to each other, they have to overcome a barrier $V_{\mathrm{SP}}$. SP is a saddle point in the full energy landscape, so the shear mode in the perpendicular direction (along the direction of two nearby AA maxima) is stable. In order to determine temperature effects on the barrier, we use SSCHA to determine the free energy at the equilibrium position and SP, using the interpolation method described in Sec. II.2. The barrier as a function of temperature can be seen in blue in Fig. 2.
@@
Figure 2: Barrier as a function of temperature. It is reduced from 1.24 meV/atom, the value at fixed nuclei (VSP in Fig. 1), to about 0.8 meV/atom, corresponding to a large reduction of about 35% due to phonons. The barrier does not change much up to about 500 K, but the value at 1000 K is about 20% smaller relative to the value at 100 K. The lines are a guide to the eye, as in the other figures of this work.
The barrier at 100 K is reduced by a substantial 35% with respect to the value at fixed nuclei. We expect intralayer modes to contribute little to the reduction of the barrier, since they should not change much when displacing the layers. On the other hand, modes involving the relative motion of the layers should be much more affected. Indeed, intralayer modes are virtually identical, so they do not contribute to a change in the barrier. Regarding the interlayer modes, the breathing mode changes little, while the shear modes do differ significantly (see Fig. S4). Thus, the shear modes are responsible for the large reduction of the barrier. This makes sense intuitively: modes that involve a relative displacement of the layers should facilitate sliding. In the fixed ion calculation, the system is exactly at SP, but in reality, the system is also exploring surrounding configurations, which have a lower energy.
Remarkably, our results show improved agreement with experiments. In a previous workNery et al. (2020), using density functional theory with an LDA functional (see Table 1), some of us obtained that the maximum shear stress in simulated stick-slip motion is 0.2 GPa, and we mentioned a good agreement with experimental values 0.14 GPa (Ref. Liu et al., 2012) –a benchmark for defect free single crystal graphiteBlomquist et al. (2019)– and 0.1 GPa (Ref. Dienwiebel et al., 2004). Assuming that the maximum shear stress decreases by 35% in line with the barrier, we now obtain 0.13 GPa, in excellent agreement with experimentsDienwiebel et al. (2004); Liu et al. (2012).
@@
The size of the graphene flakes considered in our work in principle correspond to nanometers up to possibly micrometer scales. In clean samples with micrometer sized flakes, different regions can have different stackings (AB and AC)Henck et al. (2018); Jiang et al. (2018), and a domain wall in the metastable state separates both regions. The displacement of layers in this case is more complex and is not considered in our work. We expect shear stress to be lower in this case, since the metastable state is already present. Defects and stacking faults also lead to lower friction due to incommensurabilities, so the shear stress of 0.14 GPa corresponds to the maximum expected valueBlomquist et al. (2019).
In other materials, the anharmonicities associated with the unstable modes at a saddle point should also contribute negatively to the free energy. However, the overall change of the barrier will depend on differences in the entire SSCHA phonon spectra at the minima and saddle point.
Regarding the evolution of the barrier as the temperature increases, there is a weak dependence up to about 500 K, since the barrier remains close to around 0.8 meV/atom. However, when temperature increases up to 1000 K, the barrier is further reduced by 20% relative to the value at 100 K. Thus, high temperatures should facilitate sliding between layers. In particular, it could aid the production of ABC-stacked multilayer graphene from its more common AB-stacked version through shear stress, as mentioned in the Introduction. More in general, high temperatures could also help engineer specific stacking arrangements in other layered materials.